Suppose that one worker can produce 25 cookies, two workers together can produce 75 cookies, and three workers can produce 105 cookies together. What is the marginal product of the 3rd worker?

a) 20 cookies

b) 25 cookies

c) 60 cookies

d) 30 cookies
Answer #1

Answer : [ D ] 30 cookies.

Formula :: Marginal product of labor = TP / L

= Change in Total product / Change in Labor

Calculations :

Number of Workers    Total product Marginal product
1 25 25
2 75 (75 -25)/ ( 2- 1) = 50
3 105 (105 - 75) / ( 3 - 2 ) = 30

Marginal product of the 3rd worker is 30 cookies.